# Distributional Vision-Language Alignment by Cauchy-Schwarz Divergence

### v1

@@ 3.1 Mutual Information is insufficient for alignment
$\mathcal{L}_{\text{I2T}}=-\frac{1}{N}\sum_{i=1}^{N}\log$ $\frac{\exp\left(\text{sim}(\mathbf{x}_{i},\mathbf{y}_{i})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(\mathbf{x}_{i},\mathbf{y}_{j})/\tau\right)},$
@@
$\mathcal{L}_{\text{T2I}}=-\frac{1}{N}\sum_{i=1}^{N}\log$ $\frac{\exp\left(\text{sim}(\mathbf{y}_{i},\mathbf{x}_{i})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(\mathbf{y}_{i},\mathbf{x}_{j})/\tau\right)},$
where $\text{sim}(\cdot,\cdot)$ denotes cosine similarity. $\tau$ is temperature. Critically, this formulation requires paired data $\{(x_{i},y_{i})\}$.
Although widely adopted, mutual information alone is insufficient for effective modality alignment (Liang et al., 2022). The reason is that mutual information quantifies the statistical dependence between two random variables (Cover, 1999), ensuring that the distribution $p(\mathbf{x})$ is related to $p(\mathbf{y})$. However, it does not guarantee that the distributions $p(\mathbf{x})$ and $p(\mathbf{y})$ are statistically similar or close to each other. In other words, two distributions can differ significantly or be far apart, yet exhibit strong dependence. We illustrate this issue with the following toy example.
Example 3.1.
Consider two Gaussian distributions, $p(\mathbf{x})\sim\mathcal{N}(\mu_{\mathbf{x}},\sigma_{\mathbf{x}}^{2})$ and $p(\mathbf{y})\sim\mathcal{N}(\mu_{\mathbf{y}},\sigma_{\mathbf{y}}^{2})$, with a joint distribution $p(\mathbf{x},\mathbf{y})\sim\mathcal{N}\left(\begin{pmatrix}\mu_{\mathbf{x}}\\ \mu_{\mathbf{y}}\end{pmatrix},\begin{pmatrix}\sigma_{\mathbf{x}}^{2}&\rho\sigma_{\mathbf{x}}\sigma_{\mathbf{y}}\\ \rho\sigma_{\mathbf{x}}\sigma_{\mathbf{y}}&\sigma_{\mathbf{y}}^{2}\end{pmatrix}\right)$. Here, $\mu_{\mathbf{x}}$ and $\mu_{\mathbf{y}}$ are the means of $\mathbf{x}$ and $\mathbf{y}$, $\sigma_{\mathbf{x}}^{2}$ and $\sigma_{\mathbf{y}}^{2}$ are their variances, and $\rho$ is the correlation coefficient and controls their linear dependency. When $\rho=0.99$, the two modalities are highly dependent, with high mutual information ($I=1.959$; see Fig. 2 and 2). When $\rho=0$, the modalities are independent, resulting in zero mutual information (Fig. 2). Interestingly, two distributions with the same mutual information value can either exhibit minimal statistical distance and nearly identical shapes, including similar locations, widths, and higher-order moments, as shown in Fig. 2, or have completely different shapes with distinct means (00 for $p(\mathbf{x})$ and $2$ for $p(\mathbf{y})$) and variances ($100$ for $p(\mathbf{x})$ and $0.01$ for $p(\mathbf{y})$), as illustrated in Fig. 2. Quantitatively, the former case shows a minimal KL divergence of 00, while the latter exhibits a KL divergence of nearly $5,194$. See details in Appendix A.
Example 3.1 shows that despite strong dependence and high mutual information, the representation distributions of two modalities can remain misaligned and be far from each other. This issue is also observed in the CLIP model pretrained with InfoNCE, where the vision and language representations exhibit a noticeable distributional gap, as shown in Fig. 1. This gap results in inconsistently aligned multimodal features, hindering the clear representation of shared semantics and disrupting effective mapping between modalities. Ultimately, this misalignment degrades performance in downstream tasks, including cross-modality generation.
Notably, although directly minimizing the divergence between distributions may reduce the distributional gap, it risks creating independent multimodal distributions without common semantic information (Fig. 2). Therefore, maximizing mutual information and minimizing divergence complement each other to achieve effective multimodal representation alignment.
3.2 Distributional multimodal alignment
To overcome the limitations of mutual information term alone, we propose a distributional alignment framework. Specifically, we introduce a distribution divergence minimization regularization into the optimization of multimodal alignment, defining the overall problem as maximizing mutual information while ensuring a small divergence between multimodal distributions:
$\max~{}I(\mathbf{x};\mathbf{y}),~{}\text{s.t.}~{}D(p(\mathbf{x}),p(\mathbf{y}))\leq\epsilon,$ (4)
where $\epsilon>0$ is a small constant representing the permissible divergence threshold. To address this constrained optimization, we introduce a Lagrangian multiplier $\lambda\geq 0$, reformulating it into an unconstrained problem:
$\min-I(\mathbf{x};\mathbf{y})+\lambda D(p(\mathbf{x}),p(\mathbf{y})).$ (5)
This formulation optimizes for high mutual information between paired data $(\mathbf{x},\mathbf{y})$ while reducing the divergence between the distributions $p(\mathbf{x})$ and $p(\mathbf{y})$.
Unlike parametric distributions, distributions of different real-world modalities exhibit unpredictable variability and inconsistent overlaps, meaning that $p(\mathbf{x})$ and $p(\mathbf{y})$ may follow arbitrary distributions with a small intersection.
Therefore, it is crucial to overcome these challenges to measure and optimize multimodal distribution divergence robustly. Below, we outline several key properties that an effective metric should satisfy for multimodal alignment.
@@ 3.2 Distributional multimodal alignment
Key properties for distribution align metrics:
These properties enable the divergence term in (5) to align arbitrary distributions with small support overlap, which is well-suited for large-scale multimodal applications involving deep learning.
While KL divergence and Wasserstein distance are widely used, they fail to meet these requirements. KL divergence is asymmetric, inefficient for non-Gaussian data, and unreliable for distributions with small overlap (Yu et al., 2024). Similarly, Wasserstein distance is inefficient to estimate, requiring additional learnable module (Arjovsky et al., 2017) or Sinkhorn iterations (Cuturi, 2013). As a result, these metrics are suboptimal for large-scale multimodal alignment tasks (details in Appendix B).
To satisfy these properties, we introduce CS divergence (Principe et al., 2000a, b), a symmetric and robust metric to quantify the distance between two probability density functions:
[…]
The CS divergence satisfies all the desired properties. It is a symmetric distance metric between any two probability density functions $p(\mathbf{x})$ and $p(\mathbf{y})$, satisfying $0\leq D_{\text{CS}}<\infty$, with the minimum achieved if and only if $p(\mathbf{x})=p(\mathbf{y})$. Furthermore, CS divergence can be estimated non-parametrically using a kernel density estimator (KDE) (Parzen, 1962), eliminating the need for explicit parametric assumptions about the underlying distributions. It also offers an elegant closed-form expression for mixtures of Gaussians (MoG) (Kampa et al., 2011) and infinite MoG. This provides significant flexibility in measuring distributional distance.
By incorporating the CS divergence into Eq. (5), we propose CS-Aligner, with the objective function:
$\min-I(\mathbf{x};\mathbf{y})+\lambda D_{CS}(p(\mathbf{x}),p(\mathbf{y})).$ (7)
This formulation ensures both semantic and distributional alignment, enabling robust and efficient multimodal learning across diverse real-world tasks.
3.3 CS-Aligner
We use CS-Aligner to align the pretrained multimodal models, as illustrated in Fig. 3. To enable the alignment, we detail the estimation methods for both CS divergence and mutual information in Eq. (7), providing the complete objective of CS-Aligner.
CS divergence estimation. We use KDE to estimate $D_{\text{CS}}(p(\mathbf{x});p(\mathbf{y}))$ nonparametrically. Given i.i.d. samples $\{\mathbf{x}_{i}\}_{i=1}^{M}\sim p(\mathbf{x})$ and $\{\mathbf{y}_{i}\}_{i=1}^{N}\sim p(\mathbf{y})$, the empirical CS divergence estimator is given by (Jenssen et al., 2006):
$\widehat{D}_{\text{CS}}(p(\mathbf{x});p(\mathbf{y}))=\log\Big{(}\frac{1}{M^{2}}\sum_{i,j=1}^{M}\kappa({\mathbf{x}}_{i},{\mathbf{x}}_{j})\Big{)}+$ (8)
$\log\Big{(}\frac{1}{N^{2}}\!\!\sum_{i,j=1}^{N}\!\kappa({\bf y}_{i},{\mathbf{y}}_{j})\Big{)}-2\log\Big{(}\frac{1}{MN}\!\sum_{i=1}^{M}\sum_{j=1}^{N}\!\kappa({\mathbf{x}}_{i},{\mathbf{y}}_{j})\Big{)}.$
where $\kappa$ is a kernel function such as Gaussian $\kappa_{\sigma}(\mathbf{x},\mathbf{y})=\exp(-\|\mathbf{x}-\mathbf{y}\|_{2}^{2}/2\sigma^{2})$. This estimator is symmetric, differentiable, and computationally efficient, making it suitable for multimodal alignment. Moreover, the third term in Eq. (8) ensures that $\widehat{D}_{\text{CS}}(p(\mathbf{x});p(\mathbf{y}))\to\infty$ only when $\kappa({\mathbf{x}},{\mathbf{y}})\to 0$ (i.e., when the distributions do not overlap). However, as long as there is nonzero overlap between the distributions, the estimator remains well-defined and valid.
Remark 3.3.
Connection to the prior loss ($\ell_{2}$ loss) (Patel et al., 2024). Consider the third term in Eq. (8), which involves $\kappa(\mathbf{x}_{i},\mathbf{y}_{j})$ defined by the Gaussian kernel $\kappa_{\sigma}(\mathbf{x},\mathbf{y})=\exp\!\bigl{(}-\|\mathbf{x}-\mathbf{y}\|_{2}^{2}/2\sigma^{2}\bigr{)}$. A second-order Taylor expansion yields
Mutual information estimation. The mutual information and the CS divergence serve complementary roles for alignment: mutual information captures semantic relationships by focusing on pairwise samples, while CS divergence aligns modalities at the distributional level (overall data distribution), leveraging global information.
Inspired by CLIP-based methods (Radford et al., 2021), we estimate the mutual information term $I(\mathbf{x},\mathbf{y})$ in Eq. (5)using its lower bound, optimized via InfoNCE (Eq. (3)). We specifically choose InfoNCE not only for the optimization efficiency but also for its compatibility with CS divergence in the cosine similarity space.
@@ 3.3 CS-Aligner
$\hat{D}_{\mathrm{CS}}(p(\mathbf{x});p(\mathbf{y}))$ $=-2\log\left(\frac{\langle\boldsymbol{\mu}_{x},\boldsymbol{\mu}_{y}\rangle_{\mathcal{H}}}{\|\boldsymbol{\mu}_{x}\|_{\mathcal{H}}\|\boldsymbol{\mu}_{y}\|_{\mathcal{H}}}\right)$ (10)
$=-2\log\text{sim}(\boldsymbol{\mu}_{x},\boldsymbol{\mu}_{y}),$
which evaluates the cosine similarity between distributions. Similarly, InfoNCE evaluates cosine similarity between paired samples (Eq. 3). This dual-level similarity assessment underscores the synergy between CS divergence and mutual information, offering a unified and robust framework for multimodal alignment.
Final objective function. With the exact estimation of the CS divergence in Eq. (8) and InfoNCE in Eq. (3), the final objective function of our method is formulated as:
$\mathcal{L}_{\text{CS-Aligner}}=\widehat{D}_{\text{CS}}(p(\mathbf{x});p(\mathbf{y}))+\lambda\mathcal{L}_{\text{InfoNCE}}.$ (11)
3.4 Extended alignment with unpaired data
Benefits from the distributional alignment, we further propose some novel extensions of CS-Aligner, which leverage additional information in unpaired data. While mutual information estimation requires pairwise data, the CS divergence estimator (Eq. (8)) can operate seamlessly on unpaired data without introducing additional computation. This unique capability enables CS-Aligner to extend beyond traditional pairwise multimodal alignment by incorporating additional distributional information from unpaired data or tokens. We introduce two novel directions for this extended alignment: unpaired data alignment and token alignment.
@@ 3.4 Extended alignment with unpaired data
Vision-language token alignment. We also propose a novel intra-sample distribution alignment approach between vision and language tokens. Unlike CLIP-based models (Radford et al., 2021), which align only the “CLS” tokens of vision and text representations, the method considers all vision and text tokens for a more fine-grained alignment. Specifically, each vision feature $\mathbf{x}_{i}\in\mathbb{R}^{V\times D}$ is modeled as a token distribution $p(\mathbf{x}_{i})$ containing $V$ vision tokens, while each text feature $\mathbf{y}_{i}\in\mathbb{R}^{L\times D}$ is represented as a token distribution $p(\mathbf{y}_{i})$ consisting of $L$ text tokens. $D$ denotes the feature dimension. We compute the CS divergence between the vision and text token distributions. The internal token-wise alignment loss $\mathcal{L}_{\text{token}}$ is formulated as:
@@ 3.5 Parameter-efficient multimodal alignment
We demonstrate the effectiveness of our CS-Aligner by performing vision-language alignment in a parameter-efficient manner using pretrained vision and language models, such as CLIP and large language models (LLMs) (Dubey et al., 2024). To adapt these pretrained models, we employ two widely used frameworks: adapter (Gao et al., 2024) and LoRA (Low-Rank Adaptation) (Hu et al., 2021).
Adapter alignment. We add a lightweight transformer (Vaswani, 2017) on top of the pretrained model as an adapter. The adapter projects text embeddings or image embeddings into a shared representation space and distribution.
LoRA alignment. We also explore LoRA to insert trainable low-rank matrices into the pretrained weights of the text encoder. It enables fine-grained adjustments to the representations, aligning them with the other modality distribution.
The adapter and LoRA enable efficient alignment of the multimodal large-scale pretrained models, without requiring extensive computational resources.
@@ 4 Experiments
We evaluate our method on two tasks to illustrate its vision-language alignment ability: text-to-image (T2I) generation in Section 4.1 and image-text retrieval in Section 4.2.
@@ 4.1 Text to image generation
Datasets. Following a previous T2I approach (Patel et al., 2024), we train our method on four datasets: MSCOCO (Lin et al., 2014), CC3M (Sharma et al., 2018), CC12M (Changpinyo et al., 2021), and LAION-HighResolution-5M (Schuhmann et al., 2022). MSCOCO contains 80K images paired with multiple captions. CC3M and CC12M include about 2.5M and 10M image-text pairs, respectively. LAION-HighResolution comprises 175M high-resolution pairs, from which we select 5M for training. We evaluate the aligned model on the MSCOCO 30K validation set.
Experimental setup. We build our method based on unCLIP-style approaches (e.g., DALL-E-2 (Ramesh et al., 2022), Karlo (Donghoon et al., 2022), Kandinsky (Razzhigaev et al., 2023)). These methods typically train a diffusion prior module on large-scale datasets (more than hundreds of millions data) to map text representations into the image representation space. The diffused text representations are fed into a separate decoder for image generation.
Differently, CS-Aligner trains an adapter to align text representations to image feature space on small-scale datasets, e.g., MSCOCO (0.08M), CC3M (3M), and CC12M (12M), and LAION-HighRes subset (5M). After alignment, we directly process the aligned text features using the pretrained decoder of the large-scale methods (e.g., Karlo and Kandinsky) to generate images, without additional prior modules or multiple diffusion steps. We evaluate generation quality with the FID score (Heusel et al., 2017), which measures how closely generated images match the real image distribution. This metric is particularly well-suited for evaluating modality alignment, as it directly reflects the distribution distance. Additional details can be found in Appendix C.
Baselines. Our baselines consists of both large-scale methods Karlo (Donghoon et al., 2022), Kandinsky (Razzhigaev et al., 2023) and recent small-scale alignment method Eclipse (Patel et al., 2024). Eclipse streamlines the prior module in Karlo and Kandinsky by employing an $L_{2}$ loss for T2I. For a fair comparison, we adopt the same Transformer adapter as Eclipse and only align the “CLS” tokens, highlighting the advantages of our distributional alignment.
@@
Table 2: Comparisons with Eclipse on various training data. Our method consistently performs better.
@@
Comparisons. We compare our method with both the large-scale diffusion-based methods and the small-scale alignment methods. The results are provided in Table 1. By aligning text representations to image representations on the small MSCOCO data, our method achieves superior T2I generation than the large-scale methods Karlo and Kandinsky, without any diffusion steps. CS-Aligner also outperforms Eclipse by an obvious margin using either Karlo or Kandinsky decoders. The results demonstrate the effective vision-language alignment capability of our method. Moreover, we compare CS-Aligner with Eclipse across different training datasets. As shown in Table 2, our method performs better across diverse training data (CC3M, CC12M, and LAION-HighRes-5M), underscoring the importance of the modality distribution information for robust alignment.
Figure 4: Qualitative Visualization. We present visualizations for no alignment (left), Eclipse (middle), and CS-Aligner (right). CS-Aligner achieves more realistic generations with stronger semantic consistency.
Qualitative Visualization. To further evaluate our method, we present qualitative visualizations of generated images using the Karlo decoder. As shown in Fig. 4, our aligned text representations result in more realistic images with stronger semantic consistency with the input sentence, highlighting the effectiveness of CS-Aligner in enhancing alignment.
@@
Figure 5: CS-Aligner with additional information. Our method benefits from the additional information from the multiple captions (a) and unpaired data (b).
CS-Aligner with multiple captions. It is common in real-world datasets for a single image to correspond to multiple captions (e.g., 5 captions per image in MSCOCO). Due to their pairwise alignment nature, previous methods such as InfoNCE and $\ell_{2}$-based approaches (Radford et al., 2021; Patel et al., 2024) struggle to simultaneously leverage multiple captions. In contrast, by incorporating CS divergence, our CS-Aligner enables training for alignment with single image and multiple captions. To demonstrate the benefits of multiple captions for CS-Aligner, we conducted experiments on the MSCOCO dataset by estimating the CS divergence term $\widehat{D}_{\text{CS}}$ in Eq. (11) using both single and multiple captions. As shown in Fig. 5, CS-Aligner effectively leverages the information provided by multiple captions, leading to improved vision-language alignment.
CS-Aligner with additional unpaired data. Collecting and accurately annotating paired vision-language data is both challenging and costly. Enhancing alignment with additional unpaired data offers a more flexible and scalable solution for real-world applications. However, similar to the case of multiple captions, previous methods (Radford et al., 2021; Patel et al., 2024) struggle to fully utilize unpaired data due to their reliance on pairwise alignment, whereas CS-Aligner naturally incorporates the unpaired data information by CS divergence. To demonstrate this capability, we conduct experiments on the MSCOCO dataset using the Kandinsky decoder with (1) 80K paired training samples, (2) 40K paired training samples, and (3) 40K paired training samples supplemented with 80K unpaired samples, where the unpaired samples are used to estimate the CS divergence. As shown in Fig. 5, the result with 40K paired training data is lower than 80K. However, introducing additional unpaired data obviously improves the performance, even surpassing the model trained with 80K paired samples. This demonstrates CS-Aligner’s ability to effectively leverage the distributional information of modalities for alignment.
CS-Aligner with token alignment. Beyond the unpaired data, CS-Aligner also enables token-level alignment by treating the tokens of each sample as a distribution. We evaluated the token-level extension of CS-Aligner using the Kandinsky decoder on the MSCOCO dataset. As shown in Fig. 6, incorporating token alignment further improves performance. Moreover, qualitative results indicate that token alignment enhances fine-grained details in generated images, suggesting an improved ability to capture fine-grained relationships between modalities. Additional visualizations are provided in Fig. 8 in Appendix D.2.
Figure 6: CS-Aligner with token alignment. Token alignment enhance more fine-grained vision-language alignment.

### v2

@@ 3.1 CS-Aligner: distributional multimodal alignment
$\mathcal{L}=$ $\underbrace{{\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim p_{\text{pair}}}\left[\left\|\mathbf{x}-\mathbf{y}\right\|_{2}^{\alpha}\right]}-\log\,\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x}),\mathbf{y}\sim p(\mathbf{y})}\left[\exp\left(-t\|\mathbf{x}-\mathbf{y}\|^{2}\right)\right]}_{\text{Alignment}}$ (10)
@@
$+\underbrace{\log\,\mathbb{E}_{\mathbf{x},\mathbf{x}^{\prime}\sim p(\mathbf{x})}\left[\exp\left(-t{\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}}\right)\right]}_{\text{Uniformity on \bf{x}}}+\underbrace{\log\,\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}\sim p(\mathbf{y})}\left[\exp\left(-t{\|\mathbf{y}-\mathbf{y}^{\prime}\|^{2}}\right)\right]}_{\text{Uniformity on \bf{y}}}.$
@@
For the alignment part, CS-Aligner promotes both the matching of image-text pairs and the alignment of global distributions. For uniformity, CS-Aligner encourages dispersion within each modality independently, rather than across modalities, which could otherwise conflict with the alignment objective. Thus, our method simultaneously fosters both alignment and uniformity while avoiding the potential conflicts inherent in InfoNCE.
@@
$\hat{D}_{\mathrm{CS}}(p(\mathbf{x});p(\mathbf{y}))$ $=-2\log\left(\frac{\langle\boldsymbol{\mu}_{x},\boldsymbol{\mu}_{y}\rangle_{\mathcal{H}}}{\|\boldsymbol{\mu}_{x}\|_{\mathcal{H}}\|\boldsymbol{\mu}_{y}\|_{\mathcal{H}}}\right)=-2\log\text{sim}(\boldsymbol{\mu}_{x},\boldsymbol{\mu}_{y}),$ (11)
which evaluates the cosine similarity between distributions in RKHS. Similarly, InfoNCE evaluates cosine similarity between paired samples (Eq. (2)). This dual-level similarity assessment underscores the synergy between CS divergence and mutual information, offering a unified and robust framework for multimodal alignment.
Therefore, CS divergence is compatible with InfoNCE and effectively addresses the inherent conflict between uniformity and alignment, a property not shared by other distribution distance metrics. Detailed comparisons with other metrics are provided in the Appendix D.
3.2 Extend CS-Aligner to unpaired data
Benefits from the distributional alignment, we further propose extensions of CS-Aligner, which leverage additional information in unpaired data. While the mutual information estimation (InfoNCE) part requires pairwise data, the CS divergence estimator (Eq. (8)) can operate seamlessly on unpaired data without introducing additional computation. This unique capability enables CS-Aligner to extend beyond traditional pairwise multimodal alignment by incorporating additional distributional information from unpaired data or tokens. Below, we introduce two novel directions.
@@ 3.2 Extend CS-Aligner to unpaired data
Vision-language token alignment. We propose a novel intra-sample distribution alignment approach between vision and language tokens. Unlike CLIP-based models [6] which align only the “CLS" tokens of vision and text representations, the method considers all vision and text tokens for a more fine-grained alignment. Specifically, each vision feature $\mathbf{x}_{i}\in\mathbb{R}^{V\times D}$ is modeled as a token distribution $p(\mathbf{x}_{i})$ containing $V$ vision tokens, while each text feature $\mathbf{y}_{i}\in\mathbb{R}^{L\times D}$ is represented as a token distribution $p(\mathbf{y}_{i})$ with $L$ text tokens. $D$ denotes the feature dimension. We compute CS divergence between vision and text token distributions, and obtain internal token-wise alignment loss:
@@ 3.3 Parameter-efficient multimodal alignment
We demonstrate the effectiveness of our CS-Aligner by performing vision-language alignment in a parameter-efficient manner using pretrained vision and language models, such as CLIP and large language models (LLMs) [25]. To adapt these pretrained models, we employ two widely used frameworks: adapter [13] and LoRA [26]. The adapter and LoRA enable efficient alignment of the multimodal large-scale pretrained models, without requiring extensive computational resources. The whole framework is demonstrated in Fig. 3.
Adapter alignment. We add a lightweight transformer [27] on top of the pretrained model as an adapter. The adapter projects text embeddings or image embeddings into a shared space.
LoRA alignment. We insert trainable low-rank matrices into the pretrained weights of text encoder. It enables fine-grained adjustments to the representations, aligning them with the other modality.
@@ 4 Experiments
We evaluate our method on two tasks to illustrate its vision-language alignment ability: text-to-image (T2I) generation in Section 4.1 and image-text retrieval in Section 4.2. Additionally, we provide the image-text classification results in the Appendix G.
@@ 4.1 Text to image generation
Datasets. Following a previous T2I approach [15], we train our method on four datasets: MSCOCO [16], CC3M [28], CC12M [29], and LAION-HighResolution-5M [30]. MSCOCO contains 80K images paired with multiple captions. CC3M and CC12M include about 2.5M and 10M image-text pairs, respectively. LAION-HighResolution comprises 175M high-resolution pairs, from which we select 5M for training. We evaluate the aligned model on the MSCOCO 30K validation set.
Experimental setup. We build our method based on unCLIP-style approaches (e.g., DALL-E-2 [3], Karlo [31], Kandinsky [4]). These methods typically train a diffusion prior module on large-scale datasets (hundreds of millions data) to map text representations into the image representation space. The diffused text representations are fed into a separate decoder for image generation.
Differently, CS-Aligner trains an adapter to align text representations to image feature space on small-scale datasets, e.g., MSCOCO (0.08M), CC3M (3M), and CC12M (12M), and LAION-HighRes subset (5M). After alignment, we directly process the aligned text features using the pretrained decoder of the large-scale methods (e.g., Karlo and Kandinsky) to generate images, without additional prior modules or multiple diffusion steps. We evaluate generation quality with the FID score [32], which measures how closely generated images match the real image distribution. This metric is particularly well-suited for evaluating modality alignment, as it directly reflects the distribution distance. Additional details can be found in Appendix F.
@@
Baselines. Our baselines consists of both large-scale methods Karlo [31], Kandinsky [4] and recent small-scale alignment method Eclipse [15]. Eclipse streamlines the prior module in Karlo and Kandinsky by employing an $L_{2}$ loss for T2I. For fairness, we use the same Transformer adapter as Eclipse and only align the “CLS” tokens, highlighting the advantages of our distributional alignment.
Comparisons. We compare our method with both the large-scale diffusion-based methods and the small-scale alignment methods. The results are provided in Table 1. By aligning text representations to image representations on the small MSCOCO data, our method achieves superior T2I generation than the large-scale methods, Karlo and Kandinsky, without any diffusion steps. CS-Aligner also outperforms Eclipse by an obvious margin using either Karlo or Kandinsky decoders. The results demonstrate the effective vision-language alignment capability of our method. Moreover, we compare CS-Aligner with Eclipse across different training datasets. As shown in Table 2, our method performs better across diverse training data (CC3M, CC12M, and LAION-HighRes-5M), underscoring the importance of the modality distribution information for robust alignment.
Table 2: Comparisons on various training data. Our method consistently performs better.
@@
Qualitative Visualization. To further test our method, Fig. 4(a) shows qualitative visualizations of generated images using Karlo decoder. Our aligned text representations result in more realistic images with stronger semantic consistency with the input sentence, highlighting the effectiveness of CS-Aligner in enhancing alignment. More visualizations are provided in Appendix E.1.
(a) Qualitative comparison. No alignment (left), Eclipse (middle), and CS-Aligner (right). CS-Aligner yields more realistic, semantically consistent generations.
(b) CS-Aligner with token alignment. Token alignment enhances fine-grained vision–language correspondence.
Figure 4: Qualitative visualizations.
CS-Aligner with multiple captions. It is common in real-world datasets for a single image to correspond to multiple captions (e.g., 5 captions per image in MSCOCO). Due to their pairwise alignment nature, previous methods such as InfoNCE and $\ell_{2}$-based approaches [6, 15] struggle to simultaneously leverage multiple captions. In contrast, by incorporating CS divergence, our CS-Aligner enables training for alignment with single image and multiple captions through the divergence term. To demonstrate the benefits of multiple captions for CS-Aligner, we conducted experiments on the MSCOCO dataset by estimating the CS divergence term $\widehat{D}_{\text{CS}}$ in Eq. (6) using both single and multiple captions. As shown in Fig. 5(a), CS-Aligner effectively leverages the information provided by multiple captions, leading to improved vision-language alignment.
@@
CS-Aligner with additional unpaired data. Collecting and accurately annotating paired vision-language data is both challenging and costly. Enhancing alignment with additional unpaired data offers a more flexible and scalable solution for real-world applications. However, similar to the case of multiple captions, previous methods [6, 15] struggle to fully utilize unpaired data due to their reliance on pairwise alignment, whereas CS-Aligner naturally incorporates the unpaired data information by CS divergence. To demonstrate this capability, we conduct experiments on the MSCOCO dataset using the Kandinsky decoder with (1) 80K paired training samples, (2) 40K paired training samples, and (3) 40K paired training samples supplemented with 80K unpaired samples, where the unpaired samples are used to estimate the CS divergence. As shown in Fig. 5(b), the result with 40K paired training data is lower than 80K. However, introducing additional unpaired data obviously improves the performance, even surpassing the model trained with 80K paired samples. This demonstrates CS-Aligner’s ability to effectively leverage the distributional information of modalities for alignment.
@@
Figure 5: CS-Aligner with additional information. Our method benefits from the additional information from multiple captions (a) and unpaired data (b).
CS-Aligner with token alignment. Beyond the unpaired data, CS-Aligner also enables token-level alignment by treating the tokens of each sample as a distribution. We evaluated the token-level extension of CS-Aligner using the Kandinsky decoder on the MSCOCO dataset. As shown in Fig. 4(b), incorporating token alignment further improves performance. Moreover, qualitative results indicate that token alignment enhances fine-grained details in generated images, suggesting an improved ability to capture fine-grained relationships between modalities. Additional visualizations are provided in Fig. 7 in Appendix E.2.